# Learning Model Predictive Control with Error Dynamics Regression for Autonomous Racing

### v1

@@ V Experiments & Results
  • 20th iteration lap time (ILT-20): the average lap time after running the LMPC for 20 iterations. The shorter the lap time, the better the performance of the control policy.
  • Iteration to fail (ITF): the average number of iterations before an experiment fails due to loss of control or track boundary violation. We set the maximum number of iterations in an experiment to 20 laps around the track.
@@ V-A Robustness Study on Learning Parameters
This experiment aims to study how LMPC parameter tunings affect the safety of the learning process. The motivation is that there are not many trial-and-error opportunities when tuning parameters on safety-critical hardware such as a race car, and a robust control setup should be able to handle a broad range of parameter settings without destabilizing the system. We set up simulation experiments to compare our approach against the LMPC implementation in [8] in terms of their robustness to the change in the control-rate cost (CRC) ($c_{\Delta u}$ in (4a)). Intuitively the CRC is analogous to the “learning rate” for LMPC. A higher CRC will penalize drastic changes in the control input and keep the vehicle closer to the safe set over the MPC horizon. It is therefore a key parameter when tuning the behavior of the LMPC policy.
@@
In simulation, we ran both LMPC implementations with 5 different CRCs across two orders of magnitude and recorded the lap time and average velocity at every iteration. The simulation model is formulated to be as close as possible to the 1/10th scale vehicle used in the following hardware experiment. For fairness of comparison, we purposely introduced a mismatch in the tire-road friction coefficient $\mu$ in our nominal dynamics model. Whereas a value of 0.9 is used in the simulation model, a value of 1.2 is chosen for the LMPC nominal dynamics model in (1), which is typically only seen on full scale race cars with racing slick tires. This would have induced dangerous over-confidence in the model about the surface friction, as shown in the following hardware experiments, if no dynamics learning is used.
Table I presents the key results of this experiment. We swept CRC from 1.0 to 0.01 and compared the performance of the two LMPC implementations. With a high CRC penalty (experiments 1 and 2), both [8] and ours fail to reach the optimal lap time in 20 iterations, with [8] achieving a more significant lap time reduction. However, once we decrease the CRC to 0.1, [8] starts to fail at around lap 8, whereas our method exhibits greater robustness to different settings of the CRC and can converge to the optimal lap time within 20 iterations once the CRC drops below 0.1. Figure 3 compares the per-iteration performance of both methods, showing that [8] fails with more than half of the CRC settings, while our approach was able to keep the system safe for all settings.
To understand how the accuracy of the learned model compares between the two approaches, we inspect the average norm of the difference between the learned dynamics matrix $A(\bar{z},\mathcal{D}^{j-1})$ and the actual local dynamics matrix of the simulation model, which is computed by taking the Jacobian of the true system model at the same points $\bar{z}$. Figure 2 shows the evolution of the mismatch in the dynamics matrix over laps of a simulation experiment with a CRC value of 1. We highlight that although the modeling error of full dynamics regression in [8] eventually reaches the same level as our error dynamics regression, our learning approach better approximates the actual local dynamics consistently and always only shows a small model mismatch after learning. This is especially important in the initial iterations when there is relatively scarce data to accurately fit an ATV model. We believe that this large initial model mismatch with [8] may have contributed to the failure cases with lower settings of CRC, as the LMPC is unintentionally reaching the handling limit due to exploiting a poorly fitted model before sufficient data can be collected to perform an accurate regression of the full dynamics.
Figure 2: Comparison of average learned model error per iteration for a simulation experiment.
@@
Overall, our simulation experiments suggest that LMPC with error dynamics regression exhibits greater stability and robustness to key parameter settings and works well even when scarce dynamics data is available. These properties could potentially make the system more suitable for hardware deployment in safety-critical applications.
Experiment # Control Rate Cost ILT-20 ITF
@@
1 1.0 8.5 11.8 20+ 20+
2 0.5 6.6 9.5 20+ 20+
3 0.1 - 6 9 20+
4 0.05 - 4.9 8 20+
5 0.01 - 4.8 4 20+
TABLE I: Results of the robustness study. An ITF of 20+ means that the algorithm did not fail in the 20-iteration experiment.
Figure 3: Per-iteration results of the robustness study. The left column shows [8] and the right column shows our lap time and average speed with different control-rate cost (CRC) tunings. A red cross is placed where failure occurs due to unsafe handling of the vehicle by the LMPC.
@@ V-B Hardware Experiment Comparing Learning Performance
In this section, we present results from physical experiments with the Berkeley Autonomous Race Car (BARC) platform. These robotic vehicles are based on 1/10 scale, off-the-shelf RC cars that have been heavily modified to support autonomous vehicle architecture. Their specifications are listed in Table II. LMPC computation is done off-board on a laptop computer and localization is provided by a motion capture system.
Specification BARC IAC Car
Mass 2.2 $\mathrm{kg}$ 800 $\mathrm{kg}$
Wheelbase 0.26 $\mathrm{m}$ 4 $\mathrm{m}$
Top Speed $\sim$10 $\mathrm{km}\text{/}\mathrm{h}$ $\sim$320 $\mathrm{km}\text{/}\mathrm{h}$
[…]
Max Engine Power $\sim$100 $\mathrm{W}$ $\sim$340 $\mathrm{kW}$
TABLE II: Vehicle specifications of BARC and IAC platforms used in this work.
In total, we conducted four sets of experiments. In the first two sets, the LMPC proposed in [8] is run on the BARC platform without any regressive learning of the vehicle dynamics. In the first case, we purposefully under-estimate the vehicle handling limit by reducing the tire-road friction coefficient to $\mu=0.7$ in the nominal model. In the second case, we over-estimate by setting the coefficient to $\mu=1.2$. Note that the actual tire-road friction coefficient was measured to be 0.9. The last two sets of experiments are done with the full dynamics learning from [8] and our error dynamics learning strategy, respectively. Each set of experiments consists of 10 runs under constant LMPC tuning with $\text{CRC}=0.1$, each lasting at most 20 iterations. We record the lap times and record any failure cases before the 20-iteration threshold.
Table III compares the metrics in these four sets of experiments. When the LMPC is run with $\mu=0.7$, the car is no longer able to reach the optimal lap time, presumably because the mismatched dynamics lead the controller to under-estimate the handling limit. With $\mu=1.2$, although the LMPC is still able to achieve the same lap time as our work, it fails during several trials after 17-19 iterations due to track constraint violation, exhibiting borderline unsafe behaviors. We next compare the two regressive learning strategies. Under this low CRC setting, LMPC with full dynamics regression [8] attempts to improve the lap time quickly, but eventually fails after no more than 5 iterations in all 10 trials. On the other hand, our method converges to a best lap time safely in all experiments, while exhibiting faster lap time reduction when compared to the cases of no dynamics learning, as shown in Figure 4. We note that although Figure 4 seems to suggest that simply over-estimating the vehicle handling limit could also lead to optimal lap times, it comes at the cost of a lower ITF value and is potentially more dangerous. Figure 5 visualizes the vehicle’s trajectory over iterations with our regressive learning strategy, which shows convergence towards an optimal race line.
@@
No regression ($\mu=0.7$) 6.292 20+
No regression ($\mu=1.2$) 5.42 18.9
Full Dynamics Regression ([8]) - 4.2
Error Dynamics Regression (Ours) 5.51 20+
TABLE III: Key metrics of the BARC hardware experiment. A hyphen in the ILT-20 column means that the algorithm fails before the 20th iteration. An ITF of 20+ means that the algorithm did not fail over the 20 laps.
Figure 4: Per-iteration results of the BARC hardware experiment. The four smaller plots on the left show the lap-time reduction of the 10 trials. The combined plot on the right compares the average lap time achieved. A red cross is placed where failure occurs due to unsafe handling of the vehicle by the LMPC.
@@ V-C Hardware Demonstration on a Full-Size Race Car
In this demonstration, we deployed our LMPC and error dynamics regression strategy on a full-size IndyLights race car which has been used in the Indy Autonomous Challenge [3]. The key specifications of the vehicle are listed in Table II. The race track used for this demonstration is the Putnam Park Road Course in Indiana, USA. We construct the initial dataset $\mathcal{D}^{0}$ for the vehicle by running a tracking MPC on the center line of the track at 10 $\mathrm{m}\text{/}\mathrm{s}$ for 3 laps. We note that this step could also be done with a more rudimentary controller. We then ran LMPC with error dynamics regression for 10 iterations. Figure 1 visualizes the trajectory driven by the vehicle, with a zoomed-in view to highlight how LMPC optimizes the shape of the trajectory for a hairpin turn. It can be clearly seen in Figure 1 that the vehicle starts at lower speeds closer to the center line and incrementally increases its speed through the corners and approaches the track boundary constraint.
@@ VI Conclusion
This work proposes a new error dynamics learning approach under the LMPC framework to iteratively and safely explore the boundary of vehicle handling limit in autonomos racing. We based the learning on a nominal dynamics model, which provides explanability and safety fallback for the learning process. We then derived the regression method on top to learn a local approximation of the nonlinear model mismatch. Through simulation and hardware experiments, we showed that our method exhibits greater robustness against parameter tuning and data scarcity compared with previous LMPC works, and offers new potential solutions to address the challenge of vehicle dynamics data acquisition in high-speed autonomous racing.

### v2

@@ V Experiments & Results
  • 20th-iteration lap time (ILT-20): the average lap time after running the LMPC for 20 iterations.
  • Iteration to fail (ITF): the average number of iterations before an experiment fails due to loss of control or track boundary violation.
@@ V-A Robustness Study on Learning Parameters
Figure 2: Average error of the learned model (Top) and lap times (bottom) over iterations for various settings of the bandwidth $h$. Red and blue lines correspond to the error regression and full regression cases, respectively. A cross is placed where failure occurs due to the vehicle leaving the track.
@@
This experiment aims to study how LMPC parameter tunings affect the safety of the learning process. The motivation is that there are not many trial-and-error opportunities when tuning parameters on safety-critical hardware such as a race car, and a robust control setup should be able to handle a broad range of parameter settings without destabilizing the system. We set up simulation experiments to compare our approach against the LMPC implementation in [8] in terms of their robustness to the change in the bandwidth parameter $h$ and the control-rate cost (CRC) ($c_{\Delta u}$ in (4a)). Intuitively, the bandwidth $h$ governs a trade off between the quality and quantity of the data used for model regression. Whereas a high bandwidth allows for potentially more data points to be selected, they can be far (in the sense of the norm $\|\cdot\|_{Q}$) from the linearization point. On the other hand, a low bandwidth requires that the data points be close to the linearization point, but could result in few data points being selected when data is sparse. As for the CRC, it is analogous to the “learning rate” for LMPC. A higher CRC will penalize drastic changes in the control input and keep the vehicle closer to the safe set over the MPC horizon. Both $h$ and the CRC are therefore key parameters when tuning the behavior of the LMPC policy. For all simulation studies, we use the L-shaped track shown in Figure 5 and model the rigid body with the dynamic bicycle model. The tire forces are modeled with the Pacejka tire model [24] with tire-road friction coefficient $\mu=0.9$.
In our first simulation experiment, we set $\text{CRC}=0.1$ and compare the effect of the bandwidth $h$ on the accuracy of the learned model for our error regression LMPC and the full regression LMPC from [8]. In the error regression case, we use the kinematic bicycle model as the nominal model in (1) which induces mismatch with the simulation model. To measure accuracy of the learned model, we compute the Frobenius norm of the difference between the learned dynamics matrix $A(\bar{z},\mathcal{D}^{j-1})$ and the linearized dynamics of the simulation model at each time step and record the average over each iteration. The results for $h=\{3,4,5,10\}$ are shown in Figure 2, where we may immediately observe that for high settings of $h$, both the error and full regression cases perform similarly in both the learned model accuracy and lap times. Importantly, we see that the error regression case can correct for the mismatch between the nominal and simulation models, which is shown by the green line. For low settings of $h$, it can be seen that while [8] fails after the second iteration, the error regression LMPC is able to remain stable and converge to similar performance as the high bandwidth cases. This is due to the reasons discussed in Section IV where sparsity w.r.t. $h$ can result in significant model inaccuracies in [8], but would naturally induce a fall back to the nominal model in the error regression case. Sparsity in the regression data is especially acute in the initial laps when the LMPC policy is rapidly improving vehicle performance as there can be significant differences in the data between successive laps. We note that the behavior of the error regression case is largely dependent on the accuracy of the nominal model when data is sparse and stability cannot be guaranteed for arbitrary realizations of the nominal model. However, especially in car racing, it is not unreasonable that to assume access to a fairly accurate nominal model which can be obtained through standard system identification approaches. Future work will examine the effect of distance between the nominal and true system on the stability of the error regression LMPC.
@@
In our second simulation experiment, we fix the bandwidth to $h=5$ and ran both LMPC implementations with $\text{CRC}=\{1.0,0.5,0.1,0.05,0.01\}$ and recorded the lap time at every iteration. Like the previous study, we purposely introduce mismatch in the nominal model. Though for this study, we do so through the tire-road friction coefficient. Whereas a value of 0.9 is used in the simulation model, a value of 1.2 is chosen for the LMPC nominal dynamics model in (1), which is typically only seen on full-scale race cars with racing slick tires. This would have induced dangerous over-confidence in the model about the surface friction, as shown in the following hardware experiments, if no dynamics learning is used. Table I presents the key results of this experiment, where we observe that with CRC penalties of 1.0, 0.5, and 0.1, both [8] and our approach show similar performance at the end of 20 laps, with [8] achieving slightly faster lap time reduction as evidenced in Figure 3. However, once we decrease the CRC below 0.1, we see that [8] fails to complete the 20 laps, whereas our method exhibits greater robustness to different settings of the CRC and can still converge to high performing lap times within 20 iterations.
@@
Overall, our simulation experiments suggest that LMPC with error dynamics regression exhibits greater stability and robustness to key parameter settings and works well even when scarce dynamics data are available. These properties could potentially make the system more suitable for hardware deployment in safety-critical applications.
Control Rate Cost ILT-20 ITF
@@
1.0 6.4 6.5 20+ 20+
0.5 6.0 6.2 20+ 20+
0.1 5.5 5.6 20+ 20+
0.05 - 5.2 6 20+
0.01 - 5.0 4 20+
TABLE I: Results of the simulation robustness study. A hyphen in the 20th-Iteration Lap Time (ILT-20) means that the algorithm fails before the 20th iteration. An Iteration to Fail (ITF) of 20+ means that the algorithm did not fail in the 20-iteration experiment.
Figure 3: Results from the CRC robustness study. The left and right plots show the iteration lap times over different CRC tunings for [8] and our error regression LMPC respectively. A red cross is placed where failure occurs due to the vehicle leaving the track.
@@ V-B Hardware Experiment Comparing Learning Performance
Figure 4: Per-iteration results of the BARC hardware experiment. The four smaller plots on the left show the lap-time reduction of the 10 trials. The combined plot on the right compares the average lap times achieved. A red cross is placed where failure occurs due to the vehicle leaving the track.
@@
Full Regression ([8], $h=10.0$) 6.69 20+
Error Regression (ours, $h=10.0$) 6.82 20+
Full Regression ([8], $h=3.0$) 6.53 10.5
Error Regression (ours, $h=3.0$) 6.66 19.1
TABLE II: Results of the BARC hardware experiment.
In this section, we present results from physical experiments with the Berkeley Autonomous Race Car (BARC) platform. These vehicles are based on 1/10th scale, off-the-shelf RC cars that have been modified for autonomous racing. LMPC computation is done on a laptop computer with a 2.6 GHz 9th-Gen Intel Core i7 CPU, localization is provided by a motion capture system, and communication with the vehicle is handled with ROS2. We conducted two sets of experiments to demonstrate the effect of the bandwidth $h$ on the stability and learning performance of [8] and our error regression LMPC. Each set of experiments consists of 10 runs of at most 20 laps with $\text{CRC}=0.1$. We record the lap times and record any failure cases before the 20-lap threshold.
In the first experiment, we run [8] and our error regression LMPC with $h=10$. This corresponds to the high bandwidth case in the simulation study where both policies achieved similar performance and were able to successfully complete 20 laps. The results of our experiment are shown in the top row of plots in Figure 4, where we again see that the two approaches show similar performance in lap time reduction and are able to remain stable over 20 laps. This corroborates our simulation study and shows that with proper tuning, our error regression LMPC is essentially equivalent to [8]. In the second experiment, we run [8] and our error regression LMPC with $h=3$. This corresponds to the low bandwidth case in the simulation study where we data sparsity has a destabilizing effect on [8]. The results of our experiment are shown in the bottom row of plots in Figure 4, which also supports our observations from the simulation study. In particular, we observe that when a low bandwidth setting is used, [8] failed in 5 of the 10 trials, whereas our error learning LMPC only failed in a single trial.
@@ V-C Hardware Demonstration on a Full-Size Race Car
In this demonstration, we deployed our LMPC and error dynamics regression strategy on a full-size IndyLights race car which has been used in the Indy Autonomous Challenge [3]. The race track used for this demonstration is the Putnam Park Road Course in Indiana, USA. We construct the initial dataset $\mathcal{D}^{0}$ for the vehicle by running a tracking MPC on the center line of the track at 10 $\mathrm{m}\text{/}\mathrm{s}$ for 3 laps. We note that this step could also be done with a more rudimentary controller. We then ran LMPC with error dynamics regression for 10 iterations. Figure 1 visualizes the trajectory driven by the vehicle, with a zoomed-in view to highlight how LMPC optimizes the shape of the trajectory for a hairpin turn. It can be clearly seen in Figure 1 that the vehicle starts at lower speeds closer to the center line and incrementally increases its speed through the corners and approaches the track boundary constraint.
@@ VI Conclusion
This work proposes a new error dynamics learning approach under the LMPC framework to iteratively and safely explore the boundary of vehicle handling limit in autonomous racing. We based the learning on a nominal dynamics model, which provides explanability and safety fallback for the learning process. We then derived the regression method on top to learn a local approximation of the nonlinear model mismatch. Through simulation and hardware experiments, we showed that our method exhibits greater robustness against parameter tuning and data scarcity compared with previous LMPC works, and offers new potential solutions to address the challenge of vehicle dynamics data acquisition in high-speed autonomous racing.